# Boltzmann priors for Implicit Transfer Operators

### v1

@@ 4.1 Systems
Prinz potential is a 1D potential commonly used for benchmarking MD sampling methods (Prinz et al., 2011). We set the observable functions, $a$ and $b$, eq. 3, to be the identity function for computing dynamic observables. For details, see Appendix A.5.
@@ 4.2 Boltzmann priors for training data generation
Modeling the transition density over $\Omega$ requires observing transitions across the state space. In the usual setting, only one or a few initial conditions are available when data collection starts. Here, we explore the case where a BG is available before data collection and compare it against the baseline, where only one initial condition is known.
We use a BG trained on equilibrium data of the Prinz potential to sample the initial conditions of our training trajectories. We also generate trajectories using a single starting point (crystal, $\mathbf{x}=0.75$) and compare their performance with long simulations (Figure 2). We compare the two different data generation strategies using the metric, $|\Delta\text{correlation}|$, which measures the absolute difference of the time-correlation function (dynamic observable) compared to the MD ground truth (Appendix A.7). We find that the performance of models using Boltzmann priors for data generation is superior for different lag times (Figure 2a) and number of trajectories (Figure 2b). The gap between BG and the crystal baseline increases with lag-time, and as expected decreases with the number of generated trajectories.
Figure 2: Absolute difference in correlation (lower is better) of models trained on trajectories initialized on samples from a Boltzmann Generator (BG) and a single structure (crystal, $\mathbf{x}=0.75$) for the Prinz Potential under direct sampling. The former presents superior performance for different lag times (number of trajectories = $50$) (a) and number of trajectories (b). The shaded areas correspond to $95\ \%$ confidence interval.
@@ 4.3 BoPITO efficiently samples long-term dynamics in a low-data context
We find that the BoPITO models achieve a higher accuracy for long-term dynamics compared to ITO models when data is scarce, as the equilibrium distribution is known a priori and does not need to be learned from simulation data (Figure 3). The inductive bias in BoPITO models enables them to learn long-term dynamics, even in scenarios where ITO models fail. For the Prinz Potential, we find that while ITO suffers from poor performance modeling long-term dynamics when data is scarce (Figure 3a), BoPITO models (Figure 3b) accurately capture long-term dynamics without worsening the performance on short and medium timescales. The results for Alanine Dipeptide mirror those of the Prinz potential: we find that an order of magnitude more data is needed to train an ITO (Figure 3c) to the same accuracy as a BoPITO model trained on the same data (Figure 3d). Moreover, we show in Appendix A.8 that the energies of configurations sampled from our prior Boltzmann Generator used in our experiments match closely the energies of MD samples.
Figure 3: Absolute difference in correlation (lower is better, one-step sampling) between ITO (left) and BoPITO (right) split into short, medium, and long time-scales against the number of training trajectories for the Prinz potential (top) and Alanine Dipeptide (bottom). The shaded areas correspond to a $95\ \%$ confidence interval.
@@
In practical settings, our MD data will be off-equilibrium, e.g. having sampled only one or a few of the relevant modes in the Boltzmann distribution in a given trajectory. Consequently, unless extensive data across the domain $\Omega$ can be collected, models based on such data will be biased. An alternative to collecting more simulation data is to use a multi-modal strategy where experimental data is used to fill the gaps left by simulation data and bridge to long time-scale dynamics (Salvi et al., 2016; Kolloff & Olsson, 2023). In this section we explore the potential of BoPITO interpolators to integrate dynamic observables with off-equilibrium simulation data to recover unbiased long-term dynamics.
Figure 4: Fitting an interpolator of a biased BoPITO model. Both the ITO and BoPITO models are trained on biased MD simulations of Alanine Dipeptide. We fit a BoPITO interpolator selecting the most consistent ensemble with the unbiased correlation function specified in Equation 16. Numbers with blue arrows specify the interpolation parameter, $N_{\text{int}}$.
We showcase BoPITO interpolators on Alanine Dipeptide. By removing the transitions between the modes of the Boltzmann distribution corresponding to the slowest process in the system, e.g. $\phi$ crossing 00 or $2$, we generate a biased simulation data, resembling a realistic scenario. We then train ITO and BoPITO models using these biased trajectories. For lags $>100$, we sample the BoPITO interpolator estimated by matching to an unbiased dynamic observable defined by eq. 22, allowing us to overcome a systematic error in the correlation function observed in the biased MD data and for ITO (Figure 4).
@@
5 Related work
Sampling the Boltzmann distribution
Latent space simulators and Coarse graining
Classical (fine-grained) molecular dynamics simulation can be coarse-grained beyond the Born-Oppenheimer approximation, such that atomic nuclei are merged together into ‘beads.’(Noid, 2023) This strategy, in principle enables faster simulations due to the smaller number of particles and the acceleration of kinetics caused by the coarse-graining (CG) operation (Nüske et al., 2019). CG models can be estimated to closely approximate the thermodynamics of the corresponding fine-grained system through the optimization of two equivalent variational bounds (Noid et al., 2008; Lyubartsev & Laaksonen, 1995; Ercolessi & Adams, 1994). A relaxation of this bound was recently proposed to also allow for probabilistic reconstruction of fine-grained configurations (Chennakesavalu et al., 2023). These bounds have been used extensively to build CG force-field models (Husic et al., 2020; Wang et al., 2019a; Majewski et al., 2023; Charron et al., 2023) including implicit solvation models (Chen et al., 2021; Katzberger & Riniker, 2024), using deep neural networks due to their ability to capture multibody terms (Wang et al., 2021) to ultimately accelerate the prediction of equilibrium properties of molecular systems. Similarly, the development of ‘latent space simulators’ where a learned, typically low-dimensional, latent space equipped either with a propagator (Sidky et al., 2020; Chennakesavalu et al., 2023; Wang et al., 2024), or not (Wang & Gómez-Bombarelli, 2019), is learned to enable efficient simulation. These approaches, in general aim to accelerate molecular simulations akin to BoPITO, yet due to the CG operation, the molecular dynamics (kinetics) will be accelerated, and detailed knowledge of the unbiased dynamics are needed to correct this (Nüske et al., 2019; Crommelin & Vanden-Eijnden, 2011).
Figure 5: BoPITO can incorporate unbiased dynamic observables to correct a model trained on biased data. Rows of increasing time-lag (from top to bottom). Contour plots correspond to a BoPITO interpolator. The first column shows conditional transition densities projected onto the torsion angles $\phi$ and $\psi$ (inset). The black cross indicates the initial condition. The second and third columns show marginal distributions of $\phi$ and $\psi$, respectively.
Transfer Operator surrogates
Analysis of MD data often involves building transfer operator surrogates such as Markov state models (MSM) (Schütte et al., 1998; Prinz et al., 2011; Swope et al., 2004; Husic & Pande, 2018), time-lagged independent components analysis (Molgedey & Schuster, 1994; Ziehe & Müller, 1998; Pérez-Hernández et al., 2013), Markov field models or dynamic graphical models, (Olsson & Noé, 2019; Mardt et al., 2022; Hempel et al., 2022), VAMPnets (Mardt et al., 2018; Wu & Noé, 2019), or observable operator models (Wu et al., 2015). Markov state models are time-space discrete approximations of the transfer operator and Deep Generative MSM (Wu et al., 2019) and VAMPnets (Mardt et al., 2018) learn the space discretization through deep neural networks. Dynamic graphical models or Markov field models represent a time-space discrete approximation of the transfer operator injecting a (conditional) independence assumption of molecular subsystems, when modeling the transition probability, leading to better scaling for systems with poor time-scale separation (Olsson & Noé, 2019). Apart from ITO (Schreiner et al., 2023) several other deep generative approaches for modeling the transition density of molecular dynamics have recently proposed. Timewarp where a normalizing flow is used to encode the transition density (Klein et al., 2023) with limited transferability and metropolized sampling is used to access asymptotically unbiased equilibrium samples (Hastings, 1970). Score dynamics use a DDPM to model the displacements of an initial configuration towards a time-lagged one, achieving picosecond time-steps simulation and limited transferability (Hsu et al., 2024). However, unlike in the context of MSMs (Trendelkamp-Schroer et al., 2015; Rosta & Hummer, 2014; Wu et al., 2014; 2016), there are no deep generative transition density surrogates available leveraging available information about the equilibrium distribution — BoPITO is one such method.
@@ 7 Conclusion
We introduce Boltzmann Priors for Implicit Transfer Operator Learning (BoPITO), a framework to enhance ITO learning in two ways. First, a broad sampling of configuration space is used to initialize short off-equilibrium MD simulations. Second, we parameterize the transition density as an interpolation towards a pre-trained Boltzmann Generator, improving sample efficiency by an order of magnitude. BoPITO is a principled approach to embedding prior knowledge of the stationary distribution of Markovian dynamics as an inductive bias for long-term dynamical behavior. The approach enables interpolation between models trained on off-equilibrium data and the equilibrium distribution, and we can recover accurate models of unseen dynamics when informed by unbiased observables. Consequently, BoPITO is the first method to allow for the integration of multiple sources of information into the generation of deep generative surrogates of molecular dynamics.
@@ Appendix A Appendix
A.1 Eigen decomposition of the Transfer Operator and transition density
@@
A.2 Decay of time-dependent score term
@@
$s_{N}=\mathbb{E}_{i\sim\mathbf{I},t_{\mathrm{diff}}\sim\mathcal{U}(0,T)}\left[\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{t_{i}+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t_{i}},N,t_{\text{diff}},{\bm{\theta}})\right],$ (19)
of a trained model BoPITO model of the Prinz Potential with $\hat{\lambda}=0.994$. The model remains flexible for small lags but eventually decreases to 00, sampling from the equilibrium model.
@@
A.3 The score of the transition probability and BoPITO
@@
$=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}\left(s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})f_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})+g_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})\right).$
@@
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})$ (20)
A.4 Experimental parameters
A.4.1 Boltzmann Priors for dataset generation

### v2

@@ 5.1 Systems
Prinz potential is a 1D potential commonly used for benchmarking MD sampling methods (Prinz et al., 2011). We set the observable functions, $a$ and $b$ in eq. 3, to be the identity function for computing dynamic observables. For details, see Appendix A.10.
@@
Chignolin (cln025) is a fast folding protein with $10$ residues, $166$ atoms and $93$ heavy atoms. We use molecular dynamics data previously reported by Lindorff-Larsen et al. (2011). The data is proprietary but available upon request for research purposes. The simulations were performed in explicit solvent with a $2.5$ fs time-step and the positions was saved at $200$ ps intervals. We extract all heavy atoms positions from the simulations and train models on this data. We use the fraction of native contacts Lindorff-Larsen et al. (2011) to define a dynamic observable with,
$a(\mathbf{x})=b(\mathbf{x})=\frac{\sum_{i=1}^{N_{\text{res}}}\sum_{j>i}^{N_{i}}\frac{1}{1+e^{10(d_{ij}(\mathbf{x})-d_{ij}^{*}-1)}}}{\sum_{i=1}^{N_{\text{res}}}N_{i}},$ (18)
where the first sum in the numerator iterates over all $N_{\text{res}}$ residues in the protein, while the second sum considers the $N_{i}$ native contacts of residue $i$ separated by at least seven residues in the primary sequence. Here, $d_{ij}(\mathbf{x})$ and $d_{ij}^{*}$ represent the $C_{\alpha}-C_{\alpha}$ distances of residues $i$ and $j$ in the structure $\mathbf{x}$ and the native structure, respectively. This observable quantifies the protein’s foldedness, with values ranging from 0 (unfolded) to 1 (folded). For a detailed explanation, refer to Appendix A.12.
@@ 5.2 Boltzmann priors for training data generation
Figure 2: Absolute difference in correlation with respect to long unbiased MD simulations (lower is better) of models trained on trajectories initialized on samples from a Boltzmann Generator (BG) and a single structure (crystal, $\mathbf{x}=0.75$) for the Prinz Potential under direct sampling. The former presents superior performance for different lag times (number of trajectories = $50$) (a) and number of trajectories (b). The shaded areas correspond to $95\ \%$ confidence interval.
Modeling the transition density over $\Omega$ requires observing transitions across the state space. In the usual setting, only one or a few initial conditions are available when data collection starts. Here, we explore the case where a BG is available before data collection and compare it against the baseline, where only one initial condition is known. We use a BG trained on equilibrium data of the Prinz potential to sample the initial conditions of our training trajectories. We also generate trajectories using a single starting point (crystal, $\mathbf{x}=0.75$) and compare their performance with long simulations (Figure 2). We compare the two different data generation strategies using the metric, $|\Delta\text{correlation}|$, which measures the absolute difference of the time-correlation function (dynamic observable) compared to the MD ground truth (Appendix A.13). We find that the performance of models using Boltzmann priors for data generation is superior for different lag times (Figure 2a) and number of trajectories (Figure 2b). The gap between BG and the crystal baseline increases with lag-time, and as expected decreases with the number of generated trajectories.
@@ 5.3 BoPITO efficiently samples long-term dynamics in a low-data context
We find that the BoPITO models achieve a higher accuracy for long-term dynamics compared to ITO models when data is scarce, as the equilibrium distribution is known a priori and does not need to be learned from simulation data (Figure 3). The inductive bias in BoPITO models enables them to learn long-term dynamics, even in scenarios where ITO models fail. For the Prinz Potential, we find that while ITO suffers from poor performance modeling long-term dynamics when data is scarce, BoPITO models accurately capture long-term dynamics without worsening the performance on short and medium time-scales. The results for Alanine Dipeptide and Chignolin show favorable scaling with increasing system size: we find that an order of magnitude more data is needed to train an ITO to the same accuracy as a BoPITO model trained on the same data. Moreover, we show in Appendix A.14 that the energies of configurations sampled from our prior Boltzmann Generator used in our experiments match closely the energies of MD samples.
Figure 3: Absolute difference in correlation with respect to long unbiased MD simulations (lower is better, one-step sampling) of ITO (left) and BoPITO (right) split into short, medium, and long time-scales against the number of training trajectories for the Prinz potential (top) and Alanine Dipeptide (middle) and Chignolin (bottom). The shaded areas correspond to a $95\ \%$ confidence interval.
@@
In practical settings, our MD data will be off-equilibrium, e.g. having sampled only one or a few of the relevant modes in the Boltzmann distribution $\mu(\mathbf{x})$ in a given trajectory. Consequently, unless extensive data across the domain $\Omega$ can be collected, models based on such data will be biased. An alternative to collecting more simulation data is to use a multi-modal strategy where experimental data is used to fill the gaps left by simulation data and bridge to long time-scale dynamics (Salvi et al., 2016; Kolloff & Olsson, 2023). In this section we explore the potential of BoPITO interpolators to integrate dynamic observables with off-equilibrium simulation data to recover unbiased long-term dynamics.
Figure 4: Fitting an interpolator of a biased BoPITO model for Alanine Dipeptide. Both the ITO and BoPITO models are trained on biased MD simulations of Alanine Dipeptide. We fit a BoPITO interpolator selecting the most consistent ensemble with the unbiased correlation function specified in eq. 16. Numbers with blue arrows specify the interpolation parameter, $N_{\text{int}}$.
We showcase BoPITO interpolators on Alanine Dipeptide. By removing the transitions between the modes of the Boltzmann distribution corresponding to the slowest process in the system, e.g. $\phi$ crossing 00 or $2$, we generate a biased simulation data, resembling a realistic scenario. We then train ITO and BoPITO models using these biased trajectories. For lags $>100$, we sample the BoPITO interpolator estimated by matching to an unbiased dynamic observable defined by eq. 26, allowing us to overcome a systematic error in the correlation function observed in the biased MD data and for an ITO model trained on these data (Figure 4).
@@ 6 Limitations and future work
Figure 5: BoPITO can incorporate unbiased dynamic observables to correct a model trained on biased data. Rows of increasing time-lag (from top to bottom). Contour plots correspond to a BoPITO interpolator. The first column shows conditional transition densities projected onto the torsion angles $\phi$ and $\psi$ (inset). The black cross indicates the initial condition. The second and third columns show marginal distributions of $\phi$ and $\psi$, respectively. MSM stands for a Markov State Model of the unbiased MD data.
@@ 7 Conclusion
We introduce Boltzmann Priors for Implicit transfer Operator Learning (BoPITO), a framework to enhance ITO learning in three ways. First, a broad sampling of configuration space is used to initialize short off-equilibrium MD simulations. Second, we parameterize the transition density as an interpolation towards a pre-trained Boltzmann Generator, improving sample efficiency by an order of magnitude. BoPITO is a principled approach to embedding prior knowledge of the stationary distribution of Markovian dynamics as an inductive bias for long-term dynamical behavior. Third, our approach enables interpolation between models trained on off-equilibrium data and the equilibrium distribution, and we can recover accurate models of unseen dynamics when informed by unbiased observables. Consequently, BoPITO is the first method to allow for the integration of multiple sources of information into the generation of deep generative surrogates of molecular dynamics.
Acknowledgments
The authors thank Prof. Rocío Mercado and Prof. Fredrik Johansson for their useful feedback on the manuscript. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations in this work were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre.
@@ Appendix A Appendix
A.1 Definitions
  • Configuration space: Mathematical space in which all possible states or positions of a physical system are represented. For example, in a classical MD simulation, the positions of all the atoms in the simulation.
  • Unbiased simulations: Simulations performed with standard MD, i.e. Langevin Dynamics. They are unbiased because they generate samples from the underlying transition density. Particularly, for asymptotically large simulation times, unbiased simulations sample the Boltzmann distribution $\mu(\mathbf{x})$. However, because of energy barriers in the potential energy landscape, unbiased simulations may not explore some modes of the Boltzmann distribution if the simulation time is not long enough. Therefore, generating samples with unbiased MD often leads to off-equilibrium data.
  • Off-equilibrium data: Simulation trajectories whose underlying statistics do not represent well those of the equilibrium distribution. One example is lack of ergodicity: when the simulation fails to explore some modes in the Boltzmann distribution. Even if having explored the full state space, it is possible to have off-equilibrium trajectories because of rare events, which require to be sampled “many” times to accurately represent equilibrium statistics.
  • Biased simulation: Simulations performed with modified versions of naive MD that speed up the exploration of state space. These methods are a subset of enhanced sampling and include meta-dynamics (Laio & Parrinello, 2002), replica exchange (Earl & Deem, 2005) and others (Sidler et al., 2016; Pasarkar et al., 2023). Data generated with enhanced sampling do not resemble the underlying transition density, but, often, can be re-weighted to represent the equilibrium distribution.
A.2 Eigen decomposition of the Transfer Operator and transition density
@@
A.3 Implicit Transfer Operator details
Implicit Transfer Operator (ITO) Learning (Schreiner et al., 2023) is a framework for learning surrogate models of the transition density, $p(\mathbf{x}_{N\tau}|\mathbf{x}_{0})$. ITO models are trained sampling tuples $(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i}+N_{i}\tau},N_{i})$ and training a generative model to mimic the empirical transition density at multiple time-horizons, $N\tau$. Training is done following Algorithm 1.
Algorithm 1 Training. DisExp is defined in Algorithm 4
Input: $n$ MD-trajectories; $\mathcal{X}=\{\mathbf{x}^{j}_{0},\dots,\mathbf{x}^{j}_{t_{j}}\}_{j=0}^{n}$, ITO score-model; $\hat{\epsilon}_{\theta}$, max lag; $N_{\text{max}}$
$\mathcal{X}^{\prime}=\mathrm{Concatenate}(\{\mathbf{x}^{j}_{0},\dots,\mathbf{x}^{j}_{t_{j}-N_{\text{max}}}\}_{j=0}^{n})$
while not converged do
     $\mathbf{x}_{t}\sim\text{Choice}(\mathcal{X}^{\prime})$
     $N\sim\text{DisExp}(N_{\text{max}})$
     $t_{\mathrm{diff}}\sim\text{Uniform}(0,T)$
     Take gradient step on:
     $\nabla_{\bm{\theta}}\left[\lVert\epsilon-\hat{\epsilon}_{\bm{\theta}}(\widetilde{\mathbf{x}}^{t_{\mathrm{diff}}}_{t+N\tau},\mathbf{x}_{t},N,t_{\mathrm{diff}})\rVert_{2}\right]$
end while
return $\hat{\epsilon}_{\bm{\theta}}$
Once a model is trained, it can be sampled by following Algorithm 2.
Algorithm 2 Sampling from $\hat{p}_{\bm{\theta}}(\mathbf{x}_{0},N)$
Input: initial condition $\mathbf{x}_{0}$, lag; $N$, diffusion steps; $T_{\mathrm{diff}}$, ITO score-model; $\hat{\epsilon}_{\theta}$
$\mathbf{x}_{N}^{T_{\mathrm{diff}}}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$
for $t_{\mathrm{diff}}=T_{\mathrm{diff}}\dots 1$ do
     $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{1})$
     $\mathbf{x}^{t_{\mathrm{diff}}-1}_{N}=\frac{1}{\sqrt{\alpha^{t_{\mathrm{diff}}}}}\left(\mathbf{x}_{N}^{t_{\mathrm{diff}}}-\frac{1-\alpha^{t_{\mathrm{diff}}}}{\sqrt{1-\bar{\alpha}^{t_{\mathrm{diff}}}}}\hat{\epsilon}_{\theta}(\mathbf{x}^{t_{\mathrm{diff}}}_{N},\mathbf{x}_{0},N,t_{\mathrm{diff}})\right)+\sqrt{\beta_{t}}\epsilon$
end for
return $\mathbf{x}_{N}^{0}$
Several sampling steps can be annealed to sample longer lag-times as depicted in Algorithm 3.
Algorithm 3 Ancestral sampling. Sampling from $p_{\bm{\theta}}$ is defined in Algorithm 2
Input: initial condition $\mathbf{x}_{0}$, lag $N$, ancestral steps $n$.
Allocate $\mathcal{T}\in\mathbb{R}^{(n+1)\times\mathrm{dim}(\mathbf{x}_{0})}$
$\mathcal{T}[0]=\mathbf{x}_{0}$
for $i=1\dots n$ do
     $\mathbf{x}_{i}\sim\hat{p}_{\bm{\theta}}(\mathcal{T}[i-1],N)$
     $\mathcal{T}[i]=\mathbf{x}_{i}$
end for
return $\mathcal{T}$
Algorithm 4 Sampling from DisExp
$N_{\text{log}}\sim\text{Uniform}(0,\log(N_{\text{max}}))$
Return: floor$(\exp(N_{\text{log}}))$
BoPITO uses Algorithm 1 for training and Algorithm 2 for sampling as well, but uses the score mode in eq. 14.
A.4 Decay of time-dependent score term
@@
$s_{N}=\mathbb{E}_{i\sim\mathbf{I},t_{\mathrm{diff}}\sim\mathcal{U}(0,T)}\left[|\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{t_{i}+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t_{i}},N,t_{\text{diff}},{\bm{\theta}})|\right],$ (20)
of a trained BoPITO model of the Prinz Potential with $\hat{\lambda}=0.994$. The expectation for $\mathbf{I}$ corresponding to the sampling of tuples as during training (Alg. 1). The model remains flexible for small lags but eventually decreases to 00, sampling from the equilibrium model.
@@
A.5 The score of the transition probability and BoPITO
@@
$=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}\underbrace{\left(s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})f_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})+g_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})\right)}_{s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})}.$
@@
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}}).$ (21)
For simplicity, we incorporate the structure of the transition density into the diffusion model, not only for $t_{\text{diff}}=0$ (the learned data distribution), but for all $t_{\text{diff}}>0$. We do not observe this choice to limit model expressivity in our experiments.
A.6 Fitting $\hat{\lambda}$
The hyperparameter $\hat{\lambda}$ controls the time-scale at which the BoPITO model transitions to sampling the equilibrium model. Ideally, $\hat{\lambda}$ should be similar to the eigenvalue corresponding with the slowest process in the system, $\lambda_{2}$. However, its value is not generally available and requires extensive unbiased simulation data to be accurately estimated. If $\hat{\lambda}$ is too small, the model may prematurely relax to equilibrium, limiting its ability to capture non-equilibrium dynamics. Conversely, if $\hat{\lambda}$ is too large, the benefits of the BoPITO framework could diminish. Therefore, careful tuning of $\hat{\lambda}$ can be crucial for optimal performance.
As illustrated in Figure 7, grid-search hyper-parameter tuning can select an appropriate $\hat{\lambda}$ for a single-system model. Too small values of $\hat{\lambda}$ can hinder the model’s ability to capture long-time-scale dynamics, leading to increased loss. We recommend choosing the smallest $\hat{\lambda}$ that yields a plateau in the $\hat{\lambda}$-loss curve (elbow rule), as demonstrated in Figure 7 (b) for the Prinz potential. However, practitioners should be cautious about sampling longer lags than the implied time-scale defined by $\hat{\lambda}$ if they cannot guarantee the system relaxes to equilibrium for those time-scales. Fitting curves for Alanine Dipeptide and Chignolin are shown in Figure  8.
Figure 7: Average loss against training epochs for the Prinz potential for different values of $\hat{\lambda}$ (a) and average final loss versus $\hat{\lambda}$ (b). Averages are taken w.r.t. 10 runs. Final losses are computed as the average between epochs $190$ and $200$. The greatest eigenvalue of the system under our simulation parameters is $0.994$.
Figure 8: Average final loss against different values of $\hat{\lambda}$ for Alanine Dipeptide (a) and Chignolin (b). The smallest value of $\hat{\lambda}$ that allowed numerically stable training for Chignolin was 0.9995. Orange cross symbolizes that the trainings crashed because of numerical instability reasons.
A.7 Alternative score formulations
The score in eq. 14 is principled and resembles the score of the transition density. However, in practice, we found that this score can lead to suboptimal performance in learning short time-scales if the dynamic component fails to dominate the equilibrium component for small $N$. We only observed this for Chignolin. We relate this issue to numerical limitations due to very different scale in neural network weights for different values of $N$. We remark that short time-scales are the least relevant since they can be easily sampled with MD. Still, to mitigate this effect we propose the following alternative factorization,
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=f(N)s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}}),$ (22)
where $f(N)$ is an increasing function and tends to $1$ for large $N$. One potential option that does not introduce extra hyper-parameters is
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=\frac{\hat{\lambda}^{-2N}-\hat{\lambda}^{2N}}{\hat{\lambda}^{-2N}+\hat{\lambda}^{2N}}s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}}).$ (23)
Moreover, to ensure adequate convergence to the equilibrium distribution and avoid over-fitting to unbiased MD for long time-scales, we use the following decay score for $N>N_{\text{decay}}$,
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N_{\text{decay}},t_{\text{diff}}).$ (24)
We use this version for the experiments conducted on Chignolin and set $N_{\text{decay}}=3t_{2}$, where $t_{2}$ is the unitless characteristic time of the slowest process in the system, $t_{2}=\frac{-1}{\log\lambda_{2}}$.
A.8 How BoPITO interpolators mitigate out-of-distribution effects
BoPITO interpolators mitigate out-of-distribution effects by, first, fixing $N=N_{max}$ as the argument of $s_{dyn}$. This choice ensures that none of our neural networks are evaluated outside of the training domain. Second, we alter long-lag interpolation steps with short-lag non-interpolation steps for local relaxation. Since we in practice see that that the long-lag steps occasionally generates ‘off data manifold’ states, e.g. structures with potential high energies, the local ‘relaxation’ steps projects us back onto the data manifold. In practice, just one short-lag non-interpolation step is sufficient to overcome this. Third, we use dynamic observables to fit the interpolation parameter. Without any additional source of information we don’t know how to choose $N_{int}$. However, when we use dynamic observables, we can select the interpolated ensemble which is the most consistent with experimental observables. So this helps calibrate the time-scales to stay “in distribution”.
A.9 Experimental parameters
A.9.1 Boltzmann Priors for dataset generation